# Embedding ample semigroups as $(2,1,1)$-subalgebras of inverse semigroups

### v2

@@ Abstract.
The problem of embedding an ample semigroup in an inverse semigroup as a $(2,1,1)$-type subalgebra is known to be undecidable. In this article, we investigate the problem for certain classes of ample semigroups. We also give examples of semigroups that are left (respectively, right) but not right (respectively, left) ample.
@@ 1. Motivation and preliminaries
Let $X$ be a (possibly empty) set. By a partial bijection of $X$ we mean a bijection $\beta:A\longrightarrow B$ such that $A$ and $B$ are subsets of $X$, called, respectively, the domain and codomain of $\beta$. We shall denote $A$ and $B$ by $Dom\beta$ and $Im\beta$, respectively. The set of all partial bijections of $X$ is denoted by $\mathcal{I}_{X}$. The empty map (which has empty domain and codomain) may be viewed as a special element of $\mathcal{I}_{X}$. Throughout this article, we shall be following the convention of writing maps to the right of their arguments. Also, parentheses around arguments will be dropped if there is no chance of ambiguity. Given $\beta,\gamma\in\mathcal{I}_{X}$, we define
$\beta\gamma=\beta|_{(Im\beta\cap Dom\gamma)\beta^{-1}}\circ\gamma,$ (1.1)
where $\circ$ denotes the ordinary composition of (full) maps. The binary operation defined by (1.1) turns $\mathcal{I}_{X}$ into a semigroup, called the symmetric inverse semigroup over $X$. Because the empty map acts as the zero element of $\mathcal{I}_{X}$, we shall denote it by 0. Note that for every $\beta\in\mathcal{I}_{X}$ there exists a unique partial bijection $\beta^{-1}\in\mathcal{I}_{X}$, such that $\beta\beta^{-1}$ and $\beta^{-1}\beta$ are identities on $Dom\beta$ and $Im\beta$, respectively. Particularly, $\beta\beta^{-1}\beta=\beta$ and $\beta^{-1}\beta\beta^{-1}=\beta^{-1}$.
In general, an element $a$ of a semigroup $S$ is said to be invertible if there exists a unique element $a^{-1}\in S$, called the inverse of $a$, such that $aa^{-1}a=a$ and $a^{-1}aa^{-1}=a^{-1}$. We call $S$ an inverse semigroup if all of its elements are invertible. The semigroup $\mathcal{I}_{X}$, discussed above, is a classical example of an inverse semigroup. From now on, the letter $T$, possibly with subscripts, will be reserved to only denote inverse semigroups.
A morphism $h:T_{1}\longrightarrow T_{2}$ of inverse semigroups is just a semigroup homomorphism from $T_{1}$ to $T_{2}$. Monomorphisms and isomorphisms of (inverse) semigroups are, respectively, the injective and bijective homomorphisms. The Wagner-Preston representation (see Theorems 1.1 and 1.2) asserts that every inverse semigroup can be embedded in a symmetric inverse semigroup.
@@
Given an inverse semigroup $T$ the map
@@
is a bijection for all $x\in T$. Furthermore, the map
$\rho_{{}_{T}}:T\longrightarrow\mathcal{I}_{T},\text{ given by }\,x\longmapsto\rho_{x},\,\,\forall\,x\in T,$
@@
Theorem 1.1 may also be presented in the following dual form.
@@
Let $T$ be an inverse semigroup. Then, the map
@@
is a bijection for all $x\in T$. Moreover, the map
$\lambda_{T}:T\longrightarrow\mathcal{I}_{T},\text{ defined as }\,x\longmapsto{\lambda_{x}},\,\,\forall\,x\in T,$
is a monomorphism.
@@
See, for instance, [8] Theorem 1.5.1. ∎
Clearly, for any inverse semigroup $T$ the map $\vartheta:(T)\rho_{{}_{T}}\longrightarrow(T)\lambda_{T}$, defined by $\rho_{x}\mapsto{\lambda_{x}}$, for all $x\in T$, is an isomorphism, where $\rho_{{}_{T}}$ and $\lambda_{T}$ are the representations given by Theorems 1.1 and 1.2, respectively. Also,
@@
$E(T)=\{aa^{-1}:a\in T\};$
moreover, the latter is a subsemilattice of the former. Identifying $T$ with its isomorphic copy in $\mathcal{I}_{T}$ we see that the idempotents $xx^{-1},x^{-1}x\in E(T)$ are, respectively, the identities on the ‘domain’ and ‘codomain’ of $x\in T$. This fact may be used to verify that
@@
defines a partial order on TT—interpret (1.3) as $x\leq y$ if and only if the domain of $x$ is a subset of the domain of $y$ and the latter agrees with the former on the restricted domain. It is an easy exercise to show that Condition (1.3) is equivalent to the following one:
$\forall\,x,y\in T,\,x\leq y\text{ iff }\,\exists\,e\in E(T),\text{ such that }x=ey.$ (1.4)
@@
A semigroup $S$ is called left ample if it can be embedded in an inverse semigroup $T$ such that $(x\phi)(x\phi)^{-1}\in S\phi$ for all $x\in S$, where $\phi$ is the embedding of $S$ into $T$. Identifying $S$ with its isomorphic copy in $T$, we shall henceforth write $xx^{-1}=[(x\phi^{\prime})(x\phi^{\prime})^{-1}]\phi^{\prime-1}$, where $\phi^{\prime}:S\longrightarrow(S)\phi$ is the isomorphism defined by $x\mapsto(x)\phi$ for all $x\in S$. So, we can say that a left ample semigroup comes equipped with a unary operation $x\mapsto x^{+}$, defined by $x^{+}=xx^{-1}$. Similarly, a semigroup $S$ is called right ample if there exists an inverse semigroup $T$ admitting a monomorphism $\phi:S\longrightarrow T$ such that $(x\phi)^{-1}(x\phi)\in S\phi$. We then define $x^{-1}x=\left[(x\phi^{\prime})^{-1}(x\phi^{\prime})\right]\phi^{\prime-1}\in S$ for all $x\in S$. A right ample semigroup comes endowed with the unary operation $x\mapsto x^{*}$, given by $x^{*}=x^{-1}x$. We shall refer to $T$ as an associated inverse semigroup of the (right, left) ample semigroup $S$. One may easily verify that $S$ is left (respectively, right) ample in $T$ if and only if $(S)\rho_{{}_{T}}$ is left (respectively, right) ample in $\mathcal{I}_{T}$, where $\rho_{{}_{T}}$ is the monomorphisms given by Theorem 1.1. This is also true for $(S)\lambda_{T}$, with $\lambda_{T}$ being the monomorphism given by Theorem 1.2. Note that, a left or a right ample semigroup may be viewed as an algebra of type $(2,1)$. A semigroup $S$ is called ample if it has (possibly different) associated inverse semigroups $T_{1}$ and $T_{2}$ making it, respectively, into a right and a left ample semigroup. An ample semigroup may be considered a $(2,1,1)$-type algebra. If $S$ is a left (respectively, right) ample subsemigroup of an associated inverse semigroup $T$ then we shall say that $S$ is left, (respectively, right) ample in $T$. Similarly, $S$ will be called ample in $T$ if it is both left and right ample in $T$. Obviously, every inverse semigroup is ample (in itself), and is, as such, a $(2,1,1)$-type algebra. Also, a subsemigroup $S$ of an inverse semigroup $T$ is ample in $T$ if $E(T)\subseteq S$, that is, if $S$ is full in $T$. The converse is not true; for example, $\mathbb{N}$ is ample but not full in the multiplicative monoid $\mathbb{Q}$. For more examples of (left, right) ample semigroups the reader is referred to [3] and [6] and the references contained therein.
Problem 1.3.
@@
Let $T_{1}$ and $T_{2}$ be different inverse semigroups containing isomorphic copies, say $S_{1}$ and $S_{2}$, of a non-inverse semigroup $S$. Let $S_{1}$ be left ample in $T_{1}$ and $S_{2}$ be right ample in $T_{2}$. Then, can we find an (associated) inverse semigroup $T$ making (an isomorphic copy of) $S$ into a left as well as right ample semigroup? In the language of universal algebra, this amounts to embedding $S$ in an inverse semigroup as a $(2,1,1)$-type subalgebra.
The following results imply that, in general, Problem 1.3 is undecidable even for finite ample semigroups.
@@
In Section 2, we shall consider Problem 1.3 for certain classes of (finite) semigroups. Also, to the knowledge of the authors, there exists no example of a left (respectively, right) ample semigroup that is not right (respectively, left) ample. A class of such semigroups will be provided in Section 3. We shall also consider Problem 1.3 for (left, right) rich ample semigroups that were introduced in [6]. For the convenience of the reader, their definitions have been reproduced below.
For a subsemigroup $S$ of an inverse semigroup $T$ let us define $S^{\prime}=\{s^{-1}\in T:s\in S\}$. We say that $S$ is rich left ample (respectively, rich right ample) in $T$ if for all $x,y\in S$ one has $xy^{-1}\in S\cup S^{\prime}$ (respectively, $x^{-1}y\in S\cup S^{\prime}$). Alternatively (see [6], Proposition 4.8), $S$ is rich left ample in $T$ if
  1. (1L)
    $S$ is left ample in $T$, and
  2. (2L)
    for all $x,y\in S$, $xy^{-1}=ayy^{-1}$, where $a\in S\cup S^{\prime}$.
Similarly, the conjunction of,
  1. (1R)
    $S$ is right ample in $T$, and
  2. (2R)
    for all $x,y\in S$, $x^{-1}y=x^{-1}xb$, where $b\in S\cup S^{\prime}$,
is equivalent to $S$ being rich right ample in $T$ ([6], Proposition 4.9). Obviosly, in Conditions (2L) and (2R), one may replace $yy^{-1}$ and $x^{-1}x$ by $y^{+}$ and $x^{*}$, respectively. We call $S$ rich left (right) ample if it can be embedded, as such, in an (associated) inverse semigroup. We say that $S$ is (two-sided) rich ample if it is rich left as well as rich right ample—again, the associated inverse semigroups need not be the same. If the element $a$ in Condition (2L) is determined uniquely then $S$ is said to be ultra rich left ample. Ultra rich right and ultra rich ample semigroups are defined analogously.
@@
Let a semigroup $S$ be left (respectively, right) ample in an inverse semigroup $T_{1}$, and $T_{2}$ be an inverse semigroup admitting a morphism $f:T_{1}\longrightarrow T_{2}$. Then one can easily verifiy that $(S)f$ is left (respectively, right) ample in $T_{2}$. The remark also holds for rich (left, right) ample semigroups. However, ultra rich (left, right) ample subsemigroups are only preserved by monomorphisms.
From any ample semigroup $S$ one may obtain an amalgam $(S;T_{1},T_{2};\phi_{1},\phi_{2})$, see [6] for details, where $T_{i},1\leq i\leq 2,$ are (associated) inverse semigroups admitting monomorphisms $\phi_{i}:S\longrightarrow T_{i}$ such that $S_{1}=(S)\phi_{1}$ is left ample in $T_{1}$ and $S_{2}=(S)\phi_{2}$ is right ample in $T_{2}$. Let $V_{1}$ and $V_{2}$ be the inverse hulls of $S_{1}$ and $S_{2}$ in $T_{1}$ and $T_{2}$, respectively (see [6]). Also, assume that the isomorphism $\psi=\phi_{1}^{-1}\circ\,\phi_{2}:(S)\phi_{1}\longrightarrow(S)\phi_{2}$ extends to a homomorphism, say $\psi^{\prime}$, from $V_{1}$ to $V_{2}$. Now, because every element of $V_{1}$ may be written as $x_{1}x_{2}\dots x_{n}$, where for each $i\in\{1,2,\dots,n\}$ either $x_{i}$ or $x_{i}^{-1}$ belongs to $S_{1}$, $\psi^{\prime}$ must be defined by
$(x_{1}x_{2}\dots x_{n})\psi^{\prime}=(x_{1})\overline{\psi}(x_{2})\overline{\psi}\dots(x_{n})\overline{\psi},$ (1.6)
where
[…]
Because $(V_{1})\psi^{\prime}$ is an inverse subsemigroup of $V_{2}$ containing $S_{2}$, we may further assert that $\psi^{\prime}$ is necessarily surjective. This implies by Remark 1.6 that $S_{2}=(S_{1})\psi^{\prime}$ is right as well as left ample in $V_{2}$, and hence in $T_{2}$. Furthermore, there exists a (retract) morphism $\psi^{\prime\prime}:V_{2}\longrightarrow V_{1}$ such that $\psi^{\prime\prime}\circ\psi^{\prime}$ is identity on $V_{2}$, see, for instance, [9] §5.6. Because, by a similar argument, $\psi^{\prime\prime}$ is also surjective, it follows that $\psi^{\prime\prime}=(\psi^{\prime})^{-1}$, implying that $\psi^{\prime}$ is an isomorphism. This implies that $S_{1}$ is also left as well as right ample in $V_{1}$, equivalently, in $T_{1}$. Condition (1) of the following proposition recapitulates the above discussion.
Theorem 1.7.
Let $T_{i},1\leq i\leq 2$, be inverse semigroups admitting monomorphisms $\phi_{i}:S\longrightarrow T_{i}$ from a semigroup $S$. Suppose that $(S)\phi_{1}$ is left ample in $T_{1}$ and $(S)\phi_{2}$ is right ample in $T_{2}$. Then $(S)\phi_{i}$ is ample in $T_{i}$ for each $i\in\{1,2\}$ if one of the following (equivalent) conditions is satisfied. Particularly, $(S)\phi_{i}$ is a $(2,1,1)$-type subalgebra of $T_{i}$, for each $i\in\{1,2\}$.
  1. (1)
    The isomorphism $\phi_{1}^{-1}\circ\phi_{2}:(S)\phi_{1}\longrightarrow(S)\phi_{2}$ can be extended to a homomorphism (equivalently, isomorphism) from the inverse hull of $(S)\phi_{1}$ in $T_{1}$ to the inverse hull of $(S)\phi_{2}$ in $T_{2}$.
  2. (2)
    If $x_{1}x_{2}\dots x_{n}=y_{1}y_{2}\dots y_{m}$ in $V_{1}$ then $(x_{1})\overline{\psi}(x_{2})\overline{\psi}\dots(x_{n})\overline{\psi}=(y_{1})\overline{\psi}(y_{2})\overline{\psi}\dots(y_{m})\overline{\psi}$ in $V_{2}$, where $V_{1}$, $V_{2}$ and $\overline{\psi}$ are as defined above.
  3. (3)
    The amalgam $(S;T_{1},T_{2};\phi_{1},\phi_{2})$ is weakly embeddable in an inverse semigroup (see [6] for definition of weak embedding).
Proof.
We only need to show that the three conditions are equivalent.
$(1)\Longleftrightarrow(3)$ follows from from [6] Theorem 4.2.
$(1)\Longrightarrow(2)$ is obvious.
As for the implication $(2)\Longrightarrow(1)$, observe that Condition (2) precisely says that the map $\psi^{\prime}:V_{1}\longrightarrow V_{2}$ given by (1.6) is well-defined. But then, it also follows from its very definition that $\psi^{\prime}$ is a homomorphism that extends $\phi_{1}^{-1}\circ\phi_{2}$. (That the extension is indeed an isomorphism, has already been observed in the discussion preceding the theorem.) ∎
2. Answering problem 1.3 for certain classes of ample semigroups

### v3

@@ Abstract.
In general, the problem of embedding an ample semigroup in an inverse semigroup as a $(2,1,1)$-type subalgebra is known to be undecidable. An affirmative answer to the problem is provided for certain classes of finite ample semigroups. We also give examples of semigroups that are left (respectively, right) but not right (respectively, left) ample.
@@ 1. Motivation and preliminaries
Let $X$ be a (possibly empty) set. By a partial bijection of $X$ we mean a bijection $\beta:A\longrightarrow B$ such that $A$ and $B$, called, respectively, the domain and image of $\beta$, are subsets of XX—the empty map is a partial bijection having empty domain and image. We shall denote $A$ and $B$ by $Dom\beta$ and $Im\beta$, respectively. The set of all partial bijections of $X$ is denoted by $\mathcal{I}_{X}$.
Throughout this article, the maps will be written to the right of their arguments. Accordingly, $\phi\circ\psi$, viz. a composition of full maps, will mean $\phi$ followed by $\psi$’. Also, for better readability, parentheses around the arguments will be sometimes dropped if there is no chance of ambiguity. For all $\beta,\gamma\in\mathcal{I}_{X}$, we define
$\beta\cdot\gamma=\beta|_{(Im\beta\cap Dom\gamma)\beta^{-1}}\circ\gamma\in\mathcal{I}_{X},$ (1.1)
where $\beta^{-1}:Im\beta\longrightarrow Dom\beta$ denotes the inverse of $\beta$. The binary operation (1.1) turns $\mathcal{I}_{X}$ into a semigroup, called the symmetric inverse semigroup over $X$. As usual, we shall simply write $\beta\gamma$ instead of $\beta\cdot\gamma$. The empty map, which will be denoted by 0, acts as the zero element of $\mathcal{I}_{X}$. Actually, $\mathcal{I}_{X}$ is a monoid whose identity is the (full) identity map $1_{X}$ on $X$. Note that for every $\beta\in\mathcal{I}_{X}$ there exists a unique partial bijection $\beta^{-1}\in\mathcal{I}_{X}$, namely the inverse of $\beta$, such that $\beta\beta^{-1}$ and $\beta^{-1}\beta$ are identities on $Dom\beta$ and $Im\beta$, respectively. Particularly, $\beta\beta^{-1}\beta=\beta$ and $\beta^{-1}\beta\beta^{-1}=\beta^{-1}$.
In general, an element $a$ of a semigroup $S$ is said to be invertible if there exists a unique element $a^{-1}\in S$, called the inverse of $a$, such that $aa^{-1}a=a$ and $a^{-1}aa^{-1}=a^{-1}$. We call $S$ an inverse semigroup if all of its elements are invertible. The semigroup $\mathcal{I}_{X}$, discussed above, is a typical example of an inverse semigroup. In what follows, the letter $T$, possibly with subscripts, will be reserved to only denote inverse semigroups.
A morphism $h:T_{1}\longrightarrow T_{2}$ of inverse semigroups is just a semigroup homomorphism from $T_{1}$ to $T_{2}$. Monomorphisms and isomorphisms of (inverse) semigroups are, respectively, the injective and bijective homomorphisms. The Wagner-Preston representation (see Theorem 1.1 below) asserts that every inverse semigroup can be embedded in a symmetric inverse semigroup.
@@
Let $T$ be an inverse semigroup. Then, for all $x\in T$ the map
@@
is a partial bijection of $T$. Furthermore, the map
$\rho:T\longrightarrow\mathcal{I}_{T},\text{ given by }\,x\longmapsto\rho_{x},\,\,\forall\,x\in T,$
@@
Given an inverse semigroup $T$, the map
@@
is a partial bijection of $T$ for all $x\in T$. Also, the map
$\lambda:T\longrightarrow\mathcal{I}_{T},\text{ defined as }\,x\longmapsto{\lambda_{x}},\,\,\forall\,x\in T,$
is an anti-monomorphism; i.e. $\lambda$ is injective and $(xy)\lambda=(y\lambda)(x\lambda)$, for all $x,y\in T$.
@@
Compare, for instance, with [13] Theorem 1.5.1—the ‘isomorphism’ is replaced with ‘anti-isomorphism’ because, contrary to [13], we are writing maps to the right of their arguments. ∎
Clearly, the map $\vartheta:(T)\rho\longrightarrow(T)\lambda$, defined by $\rho_{x}\longmapsto{\lambda_{x}}$, for all $x\in T$, is an anti-isomorphism, where $\rho$ and $\lambda$ are maps given by Theorems 1.1 and 1.2, respectively. One can also easily observe that
@@
$E(T)=\{xx^{-1}:x\in T\}.$
Actually, $E(T)$ is a subsemilattice of $T$. Identifying $T$ with its isomorphic copy in $\mathcal{I}_{T}$ we see that $xx^{-1},x^{-1}x\in E(T)$ are, respectively, the identities on the domain and image of $x\in T$. This fact may be used to verify that the conditoin
@@
defines a partial order on TT—one may interpret (1.2) as $x\leq y$ if and only if the domain of $x$ is a subset of the domain of $y$ and the latter agrees with the former on the restricted domain. It is an easy exercise to show that Condition (1.2) may equivalently be stated as:
$\forall\,x,y\in T,\,x\leq y\text{ iff }\,\exists\,e\in E(T),\text{ such that }x=ey.$
@@
Given an inverse semigroup $T$, let us further recall ([8], Remark 4.3) that for all $\in T, xx$$x^{-1}$ and $x^{-1}x$ are the minimum idempotents such that $x=xx^{-1}x$, i.e. for all $e,f\in E(T)$, $ex=x$ implies that $xx^{-1}\leq e$ and $xf=x$ implies that $x^{-1}x\leq f$.
A semigroup $S$ is called left ample if it can be embedded in an inverse semigroup $T$ such that $(x\phi)(x\phi)^{-1}\in S\phi$ for all $x\in S$, where $\phi$ is the embedding of $S$ into $T$. We then denote
$xx^{-1}:=[(x\phi^{\prime})(x\phi^{\prime})^{-1}]\phi^{\prime-1}\in S,\,\forall\,x\in S,$ (1.4)
where $\phi^{\prime}:S\longrightarrow S\phi$ is the isomorphism defined by $x\longmapsto x\phi$. Thus, a left ample semigroup comes equipped with the unary operation $x\longmapsto x^{+}:=xx^{-1}$. Similarly, a semigroup $S$ is called right ample if there exists an inverse semigroup $T$ admitting a monomorphism $\phi:S\longrightarrow T$ such that $(x\phi)^{-1}(x\phi)\in S\phi$ for all $x\in S$. We use the notation
$x^{-1}x:=\left[(x\phi^{\prime})^{-1}(x\phi^{\prime})\right]\phi^{\prime-1}\in S$
for all $x\in S$, where $\phi^{\prime}$ is the isomorphism considered above. A right ample semigroup is endowed with the unary operation $x\longmapsto x^{*}:=x^{-1}x$. We shall refer to $T$ as an associated inverse semigroup of the (left, right) ample semigroup $S$. If $S$ is a subsemigroup of $T$ then we shall say that $S$ is (left, right) ample in $T$. One may easily verify that $S$ is left (respectively, right) ample in $T$ if and only if $(S)\rho$ is left (respectively, right) ample in $\mathcal{I}_{T}$, where $\rho$ is the monomorphisms given by Theorem 1.1.
@@
A semigroup $S$ is called ample if it has (possibly different) associated inverse semigroups $T_{1}$ and $T_{2}$ making it, respectively, into a left and a right ample semigroup. We shall say that $S$ is ample in $T$ if it is both left and right ample in $T$. Obviously, every inverse semigroup is ample (in itself). Also, a subsemigroup $S$ of an inverse semigroup $T$ is ample in $T$ if $E(T)\subseteq S$, that is, if $S$ is full in $T$. The converse is not true; for example, $\mathbb{N}$ is ample but not full in the multiplicative monoid $\mathbb{Q}$. For more examples of (left, right) ample semigroups the reader is referred to [4] and [8], and the references contained therein. For a general background on inverse semigroups one may refer to [7] or [13],
From the universal algebraic perspective, a left ample semigroup may be considered as a $(2,1)$-type algebra $(S,\cdot,^{+})$. Similarly, a right ample semigroup can be viewed as a $(2,1)$-type algebra $(S,\cdot,^{*})$. An ample semigroup may then be treated as a $(2,1,1)$-type algebra $(S,\cdot,^{+},^{*})$. Particularly, every inverse semigroup is an algebra of the latter type; and by saying that $S$ is left (right) ample in $T$ we precisely mean that $S$ is $(2,1)$-type subalgebra of $T$. Consequently, $S$ is ample in $T$ if and only if it is a $(2,1,1)$-type subalgebra of $T$.
Problem 1.3 (The embedding problem).
@@
Let $T_{1}$ and $T_{2}$ be different inverse semigroups containing isomorphic copies, say $S_{1}$ and $S_{2}$, of a non-inverse semigroup $S$. Let $S_{1}$ be left ample in $T_{1}$ and $S_{2}$ be right ample in $T_{2}$ (whence $S$, being an ample semigroup, is a $(2,1,1)$-type algebra). Then, can we find an inverse semigroup $T$ containing an isomorphic copy of $S$ that is left as well as right ample in $T$? In terms of universal algebra, this amounts to embedding $S$ in an inverse semigroup as a $(2,1,1)$-type subalgebra.
The following results imply that, in general, Problem 1.3 is undecidable.
@@
In the next section we answer Problem 1.3 for certain classes of finite semigroups. Particularly, we prove that a finite ample semigroups $S$ with central idempotents is embeddable in $\mathcal{I}_{S}$ as a $(2,1,1)$-subalgebra. Then, in the last section, we provide examples of semigroups that are left (respectively, right) ample but not right (respectively, left) ample. (To the knowledge of the authors, no such example exists in the literature.)
@@
Let a semigroup $S$ be left (respectively, right) ample in an inverse semigroup $T_{1}$, and $T_{2}$ be an inverse semigroup admitting a homomorphism $f:T_{1}\longrightarrow T_{2}$. Then one can easily verify that $(S)f$ is left (respectively, right) ample in $T_{2}$.
2. Answering the embedding problem for some finite ample semigroups